A block weighing 98.0 N falls 64.0 m.
(a) What is the potential energy of the block at 64.0 m?
(b) What speed does the block have as it strikes the ground?
1
Expert's answer
2013-01-15T08:38:15-0500
(a) Potential energy isE=F*h = 98.0 N *64.0 m = 6272 J
(b)velocity we can find from energy conservation
E=mv*2/2= 6272J
wherem is mass, m=F/g = 98/9.8 = 10kg
v= sqrt(2E/m) = sqrt(2*6272J/10) = 35,4 m/s